OpenStudy (anonymous):

olve for x: -4|x + 5| = -16

OpenStudy (anonymous):

there are 2 different answers i have no idea what to do

OpenStudy (chosenmatt):

what do you think?

OpenStudy (calculusfunctions):

First divide both sides by -4 to isolate the absolute value expression.

OpenStudy (anonymous):

so -1?

OpenStudy (calculusfunctions):

Now if |f(x)| = c where c is a positive constant, then the solution is f(x) = c OR f(x) = -c

OpenStudy (calculusfunctions):

No!! If you divide both sides of the given equation by -4, then |x + 5| = 4. Where do you get -1?? lol

OpenStudy (anonymous):

i subtracted 5 from both sides after that and got x =-1

OpenStudy (calculusfunctions):

OH yes, one of the solutions -1 but that's not the only solution. What's the complete final answer?

OpenStudy (anonymous):

x=9? as the other solution

OpenStudy (calculusfunctions):

|f(x)| = c where c is a positive constant, f(x) = c OR f(x) = -c If you understand that, then try again. Show me your steps!!

OpenStudy (calculusfunctions):

f(x) = x + 5, and c = 4. Hence if |x + 5| = 4, then x + 5 = 4 and x + 5 = -4. Thus what are the solutions??

OpenStudy (jhannybean):

-4|x + 5| = -16 Isolate `|x+5|` |x+5| = 4 Understand that `|x+5|` means `x+5` and `-(x+5)` x + 5 = 4 -x - 5 = 4 x = 4 - 5 -x = 4 + 5 x = -(4 + 5)

OpenStudy (jhannybean):

Solve for both and you will get 2 values of x.

OpenStudy (calculusfunctions):

Don't type out the complete solution @Jhannybean . He needs to make the effort to show his work. How is going to learn?